Combustion of hydrocarbons such as octane (CH) produces carbon dioxide, a greenhouse gas. Greenhouse gases in the Earths a
Answer #1

1 262 H18 (1) + 2502(8) 16 C02 (9) +18H0 (9) 0.250 kg=0.250X/ 250 mol n moles = 114.23 250 (n) - 250 g octane (mm = 114.2391